# An ordering for the strength of functional dependence

### v1

@@ Appendix B Proofs
’(i) $\Longrightarrow$ (ii)’: For $x_{n}\uparrow x$ and $y_{n}\uparrow y$, continuity of measures yields
$P(Y\geq y,X\geq x)$ $=\lim_{n\to\infty}P(Y>y_{n},X>x_{n})$
$\leq\lim_{n\to\infty}P(Y^{\prime}>y_{n},X^{\prime}>x_{n})=P(Y^{\prime}\geq y,X^{\prime}\geq x).$
’(ii) $\Longrightarrow$ (i)’: For $x_{n}\downarrow x$ and $y_{n}\downarrow y$, continuity of measures yields
$P(Y>y,X>x)$ $=\lim_{n\to\infty}P(Y\geq y_{n},X\geq x_{n})$
$\leq\lim_{n\to\infty}P(Y^{\prime}\geq y_{n},X^{\prime}\geq x_{n})=P(Y^{\prime}>y,X^{\prime}>x).$
’(i) $\Longrightarrow$ (iii)’: The definition of the concordance order implies $P(X\leq x)\leq P(X^{\prime}\leq x)$ and $P(X>x)\leq P(X^{\prime}>x)$ for all $x$. This yields $P(X\leq x)\leq P(X^{\prime}\leq x)$ and $P(X\leq x)\geq P(X^{\prime}\leq x)$ for all $x$ and thus $X\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X^{\prime}$. Similarly, it follows that $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$.
’(iii) $\Longrightarrow$ (i)’: By the inclusion-exclusion principle, we have
$P(X>x,Y>y)$ $=1-P(X\leq x)-P(Y\leq y)+P(X\leq x,Y\leq y)$
$\leq 1-P(X^{\prime}\leq x)-P(Y^{\prime}\leq y)+P(X^{\prime}\leq x,Y^{\prime}\leq y)$
$=P(X^{\prime}>x,Y^{\prime}>y).$
’(i) $\Longrightarrow$ (iv)’: Similar as above, we obtain $X\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X^{\prime}$ and $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$. Further, for $x_{n}\uparrow x$ and $y_{n}\uparrow y$, it follows that
$P(Y\geq y,X\geq x)$ $=\lim_{n\to\infty}P(Y>y_{n},X>x_{n})$
$\leq\lim_{n\to\infty}P(Y^{\prime}>y_{n},X^{\prime}>x_{n})=P(Y^{\prime}\geq y,X^{\prime}\geq x).$
’(iv) $\Longrightarrow$ (i)’: Follows similar as above using an approximation of measures and the inclusion-exclusion principle. $\blacksquare$
@@
Assume (i). Denote by $N_{1}\subset[0,1]$ the exceptional $\lambda$-null set such that $P(Y\leq q_{Y}(v),U\leq u)\leq P(Y^{\prime}\leq q_{Y^{\prime}}(v),U\leq u)$ for all $v\in(0,1)\setminus N_{1}$ and $u\in(0,1)$. Noting that the marginal constraint implies $P(Y\leq q_{Y}(v))=P(Y^{\prime}\leq q_{Y^{\prime}}(v))$ for all $v\in[0,1]$ outside a $\lambda$-null set $N_{2}$ (see Lemma A.2), we obtain from the inclusion-exclusion principle that
$P(Y>q_{Y}(v),U>u)\leq P(Y^{\prime}>q_{Y^{\prime}}(v),U>u)$ (40)
@@
for $\lambda$-almost all $v\in(0,1)$ and for all $u\in(0,1)$. Then, using the identities $\{Y\geq q_{Y}(v)\}=\{F_{Y}(Y)\geq v\}$ and $\{Y^{\prime}\geq q_{Y^{\prime}}(v)\}=\{F_{Y^{\prime}}(Y^{\prime})\geq v\}$ (see Proposition A.1(i)), we obtain
$P(F_{Y}(Y)\geq v,U\geq u)\leq P(F_{Y^{\prime}}(Y^{\prime})\geq v,U\geq u)$ (42)
@@
for $\lambda$-almost all $v\in(0,1)$ and for all $u\in(0,1)$. Finally, continuity of measures from above yields statement (ii). To prove (41), we consider two cases. For $v\in\mathrm{int}(\overline{\mathsf{Ran}(F_{Y})})\setminus N$, choose $v_{n}\in\mathrm{int}(\overline{\mathsf{Ran}(F_{Y})})\setminus N$ with $v_{n}\uparrow v$. Then we have $q_{Y}(v_{n})\uparrow q_{Y}(v)$ and, using the marginal constraint, we also obtain $q_{Y^{\prime}}(v_{n})\uparrow q_{Y^{\prime}}(v)$. Hence, by continuity of measures from above, we obtain with $u_{n}\uparrow u$ that
$\begin{split}P(Y\geq q_{Y}(v),U\geq u)&=\lim_{n\to\infty}P(Y>q_{Y}(v_{n}),U>u_{n})\\ &\leq\lim_{n\to\infty}P(Y^{\prime}>q_{Y^{\prime}}(v_{n}),U>u_{n})=P(Y^{\prime}\geq q_{Y^{\prime}}(v),U^{\prime}\geq u),\end{split}$ (43)
@@
where the inequality follows from (40). For the second case, let $v\in(0,1)\setminus(\overline{\mathsf{Ran}(F_{Y})})\cup N)$ and define $v_{0}:=F_{Y}^{-}(q_{Y}(v))$. Note that, by the marginal constraint and by left-continuity of the transformation $\iota_{F_{Y}}^{-}=F_{Y}^{-}\circ F_{Y}^{-1}$ in Proposition A.1(iii), we also have $v_{0}=F_{Y^{\prime}}^{-}(q_{Y^{\prime}}(v))$. Now, choose $v_{n}\in[0,1]\setminus N$ with $v_{n}\uparrow v_{0}$. Then
$\{Y\geq q_{Y}(v)\}=\{Y\geq q_{Y}(v_{0})\}=\bigcap_{n\in\mathbb{N}}\{Y>q_{Y}(v_{n})\}\cap\{Y\notin[q_{Y}(v_{0}),q_{Y}(v))\}.$ (44)
@@
Using that $\{Y\in[q_{Y}(v_{0}),q_{Y}(v))\}$ is a $P$-null set and using a similar reasoning for $\{Y^{\prime}\geq q_{Y^{\prime}}(v)\}$, we obtain (43) also in the second case. This proves (41) and thus statement (ii).
@@
For the reverse direction, assume that (42) holds for all $(u,v)\in(0,1)^{2}$. Using $\{Y\geq q_{Y}(v)\}=\{F_{Y}(Y)\geq v\}$ and $\{Y^{\prime}\geq q_{Y^{\prime}}(v)\}=\{F_{Y^{\prime}}(Y^{\prime})\geq v\}$, we obtain (41) for all $(u,v)\in(0,1)^{2}$. We aim to show (40) for $\lambda$-almost all $v\in(0,1)$ and for all $u\in[0,1]$, which implies statement (i). To this end, consider first the case where $v\in\mathrm{int}(\overline{\mathsf{Ran}(F_{Y})})$. Choose $v_{n}\downarrow v$ and $u_{n}\downarrow u$. Then we have $\bigcap_{n\in\mathbb{N}}\{Y\geq q_{Y}(v_{n})\}=\{Y>q_{Y}(v)\}\setminus\{Y\in(q_{Y}(v),q_{Y}^{+}(v)]\}$ for $q_{Y}^{+}(v)=\sup\{y\mid F_{Y}(y)\leq v\}$. Since $\{Y\in(q_{Y}(v),q_{Y}^{+}(v)]\}$ is a $P$-null set and due to a similar reasoning for $Y^{\prime}$ (using the marginal constraint), we obtain (40). For the second case, let $v\in(0,1)\setminus\overline{\mathsf{Ran}(F_{Y})}$ and define $v_{0}:=F_{Y}(q_{Y}(v))=F_{Y^{\prime}}(q_{Y^{\prime}}(v))$. Then we have $\bigcap_{n\in\mathbb{N}}\{Y\geq q_{Y}(v_{n})\}=\{Y>q_{Y}(v_{0})\}\setminus N_{v_{0}}$ for some $Y$-null set $N_{v_{0}}$. Using a similar reasoning for an approximation of $\{Y^{\prime}>q_{Y^{\prime}}(v)\}$ from above, we obtain (40) from (41). $\blacksquare$
@@
First, recall that a comparison in the concordance order implies identical marginal distributions; see Proposition B.2. Thus (ii) implies
@@
$F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}F_{Y^{\prime}}\circ q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V)\,.$ (45)
@@
Second, $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ implies the marginal constraint $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$ which is equivalent to
$F_{Y}\circ q_{Y}(v)=F_{Y^{\prime}}\circ q_{Y^{\prime}}(v)\quad\text{for }\lambda\text{-almost all }v\in(0,1);$ (46)
@@
see Lemma A.2. Since $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}q_{Y|\mathbb{X}}^{\uparrow U}(V)$ and $Y^{\prime}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V)$ due to Proposition 2.7(i), we obtain that (45) and (46) are equivalent.
@@
Now, to show ’(i) $\Longrightarrow$ (ii)’, assume that $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\,.$ Then we obtain
@@
$\begin{split}P(q_{Y|\mathbb{X}}^{\uparrow U}(V)\leq q_{Y}(v),U\leq u)&=\int_{0}^{u}P\left(q^{\uparrow U}_{Y|\mathbb{X}}(V)\leq q_{Y}(v)\mid U=s\right)\mathsf{\,d}s\\ &=\int_{0}^{u}P\left(q^{\uparrow s}_{Y|\mathbb{X}}(V)\leq q_{Y}(v)\right)\mathsf{\,d}s\\ &=\int_{0}^{u}P\left(V\leq F_{s}(q_{Y}(v))\right)\mathsf{\,d}s\\ &=\int_{0}^{u}(\eta_{Y|\mathbb{X}}^{v})^{*}(s)\mathsf{\,d}s\\ &\leq\int_{0}^{u}(\eta_{Y^{\prime}|\mathbb{X}^{\prime}}^{v})^{*}(s)\mathsf{\,d}s\\ &=\cdots=P(q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V)\leq q_{Y^{\prime}}(v),U\leq u)\end{split}$ (47)
@@
for $\lambda$-almost all $v\in(0,1)$ and for all $u\in[0,1]$ with equality for $u=1,$ where we use for the inequality the characterization of $\preccurlyeq_{ccx}$ by the Schur order; see Theorem 2.2. The first equality is due to the disintegration theorem, the second equality holds by independence of $U$ and $V.$ For the third equality, we apply Proposition A.1(i) using that $F_{s}$ is a distribution function with associated quantile function $q_{Y|\mathbb{X}}^{\uparrow s}$; see (10). The fourth equality follows with (9) noting that $V$ is uniform on $(0,1)$.
@@
Using again $q_{Y|\mathbb{X}}^{\uparrow U}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y$ and $q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$ and using the marginal constraint, we apply Lemma B.3 to prove that (47) implies
$P\big(F_{Y}(q_{Y|\mathbb{X}}^{\uparrow U}(V))\geq v,U\geq u\big)\leq P\big(F_{Y^{\prime}}(q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V))\geq v,U\geq u\big)$ (48)
for all $(u,v)\in(0,1)^{2}$. Now, statement (ii) follows from (48) by Proposition B.2 where we use the identity in (45).
@@
To prove the reverse direction, assume that $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)\leq_{c}(F_{Y^{\prime}}\circ q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V),U)$. Applying Proposition B.2, we obtain (45) and the inequality (48) for all $(u,v)\in[0,1]^{2}$. Using that (45) is equivalent to (46) which implies the marginal constraint, we obtain from (48) by Lemma B.3 that
$P\big(q_{Y|\mathbb{X}}^{\uparrow U}(V)\leq q_{Y}(v),U\leq u\big)\leq P\big(q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V)\leq q_{Y^{\prime}}(v),U\leq u\big)$ (49)
@@
for $\lambda$-almost all $v\in(0,1)$ and for all $u\in(0,1)$, where $U$ and $V$ are independent and uniform on $(0,1)$. By a similar reasoning as in (47), it follows that
$\int_{0}^{u}(\eta_{Y|\mathbb{X}}^{v})^{*}(s)\mathsf{\,d}s$ $=P\big(q_{Y|\mathbb{X}}^{\uparrow U}(V)\leq q_{Y}(v),U\leq u\big)$
$\leq P\big(q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V)\leq q_{Y^{\prime}}(v),U\leq u\big)=\int_{0}^{u}(\eta_{Y^{\prime}|\mathbb{X}^{\prime}}^{v})^{*}(s)\mathsf{\,d}s$

### v2

@@ B Proofs
The proof follows from standard arguments like continuity of measures and the inclusion-exclusion principle. $\blacksquare$
@@
Assume i. Denote by $N_{1}\subset[0,1]$ the exceptional $\lambda$-null set such that $P(Y\leq q_{Y}(v),U\leq u)\leq P(Y^{\prime}\leq q_{Y^{\prime}}(v),U\leq u)$ for all $v\in(0,1)\setminus N_{1}$ and $u\in(0,1)$. Noting that the marginal constraint implies $P(Y\leq q_{Y}(v))=P(Y^{\prime}\leq q_{Y^{\prime}}(v))$ for all $v\in[0,1]$ outside a $\lambda$-null set $N_{2}$ (see Lemma A.2), we obtain from the inclusion-exclusion principle that
$P(Y>q_{Y}(v),U>u)\leq P(Y^{\prime}>q_{Y^{\prime}}(v),U>u)$ (B.5)
@@
for $\lambda$-almost all $v\in(0,1)$ and for all $u\in(0,1)$. Then, using the identities $\{Y\geq q_{Y}(v)\}=\{F_{Y}(Y)\geq v\}$ and $\{Y^{\prime}\geq q_{Y^{\prime}}(v)\}=\{F_{Y^{\prime}}(Y^{\prime})\geq v\}$ (see Proposition A.1i), we obtain
$P(F_{Y}(Y)\geq v,U\geq u)\leq P(F_{Y^{\prime}}(Y^{\prime})\geq v,U\geq u)$ (B.7)
@@
for $\lambda$-almost all $v\in(0,1)$ and for all $u\in(0,1)$. Finally, continuity of measures from above yields statement ii. To prove (B.6), we consider two cases. For $v\in\mathrm{int}(\overline{\mathsf{Ran}(F_{Y})})\setminus N$, choose $v_{n}\in\mathrm{int}(\overline{\mathsf{Ran}(F_{Y})})\setminus N$ with $v_{n}\uparrow v$. Then we have $q_{Y}(v_{n})\uparrow q_{Y}(v)$ and, using the marginal constraint, we also obtain $q_{Y^{\prime}}(v_{n})\uparrow q_{Y^{\prime}}(v)$. Hence, by continuity of measures from above, we obtain with $u_{n}\uparrow u$ that
$\begin{split}P(Y\geq q_{Y}(v),U\geq u)&=\lim_{n\to\infty}P(Y>q_{Y}(v_{n}),U>u_{n})\\ &\leq\lim_{n\to\infty}P(Y^{\prime}>q_{Y^{\prime}}(v_{n}),U>u_{n})=P(Y^{\prime}\geq q_{Y^{\prime}}(v),U^{\prime}\geq u),\end{split}$ (B.8)
@@
where the inequality follows from (B.5). For the second case, let $v\in(0,1)\setminus(\overline{\mathsf{Ran}(F_{Y})})\cup N)$ and define $v_{0}:=F_{Y}^{-}(q_{Y}(v))$. Note that, by the marginal constraint and by left-continuity of the transformation $\iota_{F_{Y}}^{-}=F_{Y}^{-}\circ F_{Y}^{-1}$ in Proposition A.1iv, we also have $v_{0}=F_{Y^{\prime}}^{-}(q_{Y^{\prime}}(v))$. Now, choose $v_{n}\in[0,1]\setminus N$ with $v_{n}\uparrow v_{0}$. Then
$\{Y\geq q_{Y}(v)\}=\{Y\geq q_{Y}(v_{0})\}=\bigcap_{n\in\mathbb{N}}\{Y>q_{Y}(v_{n})\}\cap\{Y\notin[q_{Y}(v_{0}),q_{Y}(v))\}.$ (B.9)
@@
Using that $\{Y\in[q_{Y}(v_{0}),q_{Y}(v))\}$ is a $P$-null set and using a similar reasoning for $\{Y^{\prime}\geq q_{Y^{\prime}}(v)\}$, we obtain (B.8) also in the second case. This proves (B.6) and thus statement ii.
@@
For the reverse direction, assume that (B.7) holds for all $(u,v)\in(0,1)^{2}$. Using $\{Y\geq q_{Y}(v)\}=\{F_{Y}(Y)\geq v\}$ and $\{Y^{\prime}\geq q_{Y^{\prime}}(v)\}=\{F_{Y^{\prime}}(Y^{\prime})\geq v\}$, we obtain (B.6) for all $(u,v)\in(0,1)^{2}$. We aim to show (B.5) for $\lambda$-almost all $v\in(0,1)$ and for all $u\in[0,1]$, which implies statement i. To this end, consider first the case where $v\in\mathrm{int}(\overline{\mathsf{Ran}(F_{Y})})$. Choose $v_{n}\downarrow v$ and $u_{n}\downarrow u$. Then we have $\bigcap_{n\in\mathbb{N}}\{Y\geq q_{Y}(v_{n})\}=\{Y>q_{Y}(v)\}\setminus\{Y\in(q_{Y}(v),q_{Y}^{+}(v)]\}$ for $q_{Y}^{+}(v)=\sup\{y\mid F_{Y}(y)\leq v\}$. Since $\{Y\in(q_{Y}(v),q_{Y}^{+}(v)]\}$ is a $P$-null set and due to a similar reasoning for $Y^{\prime}$ (using the marginal constraint), we obtain (B.5). For the second case, let $v\in(0,1)\setminus\overline{\mathsf{Ran}(F_{Y})}$ and define $v_{0}:=F_{Y}(q_{Y}(v))=F_{Y^{\prime}}(q_{Y^{\prime}}(v))$. Then we have $\bigcap_{n\in\mathbb{N}}\{Y\geq q_{Y}(v_{n})\}=\{Y>q_{Y}(v_{0})\}\setminus N_{v_{0}}$ for some $Y$-null set $N_{v_{0}}$. Using a similar reasoning for an approximation of $\{Y^{\prime}>q_{Y^{\prime}}(v)\}$ from above, we obtain (B.5) from (B.6). $\blacksquare$
The following lemma addresses distributional properties of quantile-based constructions.
Lemma B.4
Let $(W,\mathbb{X})$ be a $(1+p)$-dimensional random vector. Let $V$ be uniform on $(0,1)$ and independent from $\mathbb{X}$. Then we have
  1. (i)
    $F_{W|\mathbb{X}}^{-1}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}W$,
  2. (ii)
    $P(F_{W|\mathbb{X}}^{-1}(V)\leq y,V\leq y^{\prime})=\int_{\mathbb{R}^{p}}\min\{F_{W|\mathbb{X}=\mathbb{x}}(y),y^{\prime}\}\mathsf{\,d}P^{\mathbb{X}}(\mathbb{x})$ for all $y,y^{\prime}\in\mathbb{R}$.
The first statement is a special case of (A.3). For the second statement, we obtain from disintegration
$P\big(F_{W|\mathbb{X}}^{-1}(V)\leq y,V\leq y^{\prime}\big)$ $=\int_{\mathbb{R}^{p}}P\big(F_{W|\mathbb{X}=\mathbb{x}}^{-1}(V)\leq y,V\leq y^{\prime})\mathsf{\,d}P^{\mathbb{X}}(\mathbb{x})$
$=\int_{\mathbb{R}^{p}}\min\{F_{W|\mathbb{X}=\mathbb{x}}(y),y^{\prime}\}\mathsf{\,d}P^{\mathbb{X}}(\mathbb{x}),$
where we use independence of $\mathbb{X}$ and $V$ for the first equality. The second equality follows with the transformation in Proposition A.1i and the fact that $V$ is uniform on $(0,1)$. $\blacksquare$
@@
First, recall that a comparison in the concordance order implies identical marginal distributions; see Proposition B.2. Thus ii and iii imply
@@
$F_{Y}\circ F_{Y|\mathbb{X}}^{-1}(V)$ $\stackrel{{\scriptstyle\mathrm{d}}}{{=}}F_{Y^{\prime}}\circ F_{Y^{\prime}|\mathbb{X}^{\prime}}^{-1}(V)\quad\text{and}$ (B.10)
$F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V)$ $\stackrel{{\scriptstyle\mathrm{d}}}{{=}}F_{Y^{\prime}}\circ q_{Y^{\prime};\mathbb{X}^{\prime}}^{\uparrow U}(V)\,$, (B.11)
@@
respectively. Second, $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ implies the marginal constraint $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$ which is equivalent to
$F_{Y}\circ q_{Y}(v)=F_{Y^{\prime}}\circ q_{Y^{\prime}}(v)\quad\text{for }\lambda\text{-almost all }v\in(0,1);$ (B.12)
@@
see Lemma A.2. Since $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}q_{Y;\mathbb{X}}^{\uparrow U}(V)$ and $Y^{\prime}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}q_{Y^{\prime};\mathbb{X}^{\prime}}^{\uparrow U}(V)$ due to Theorem 2.7i and $F_{Y|\mathbb{X}}^{-1}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y$ and $F_{Y^{\prime}|\mathbb{X}^{\prime}}^{-1}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$ due to Lemma B.4, we obtain that (B.10), (B.11) and (B.12) are equivalent.
@@
To show ’i $\Longrightarrow$ ii’, assume that $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$. Then, by Proposition A.4, there exists a Borel set $A\subset(0,1)$ with $\lambda(A)=1$ such that $P(Y\leq q_{Y}(v)\mid\mathbb{X})\leq_{cx}P(Y^{\prime}\leq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})$ for all $v\in A$. Since $x\mapsto-\min\{x,\alpha\}$ is convex for all $\alpha\in\mathbb{R}$, we obtain for $v\in A$ and $v^{\prime}\in(0,1)$ that
$P(F_{Y|\mathbb{X}}^{-1}(V)\leq q_{Y}(v),V\leq v^{\prime})$ $=\int_{\mathbb{R}^{p}}\min\{F_{Y|\mathbb{X}=\mathbb{x}}(q_{Y}(v)),v^{\prime}\}\mathsf{\,d}P^{\mathbb{X}}(\mathbb{x})$
$\geq\int_{\mathbb{R}^{p^{\prime}}}\min\{F_{Y^{\prime}|\mathbb{X}^{\prime}=\mathbb{x}}(q_{Y}(v)),v^{\prime}\}\mathsf{\,d}P^{\mathbb{X}^{\prime}}(\mathbb{x})$ (B.13)
$=P(F_{Y^{\prime}|\mathbb{X}^{\prime}}^{-1}(V)\leq q_{Y}(v),V\leq v^{\prime}),$ (B.14)
where the equalities are due to Lemma B.4ii. Note that $F_{Y|\mathbb{X}}^{-1}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y$; see Lemma B.4i. Then, applying Lemma B.3 and Proposition B.2, yields $(F_{Y}\circ F_{Y|\mathbb{X}}^{-1}(V),V)\geq_{c}(F_{Y}\circ F_{Y^{\prime}|\mathbb{X}^{\prime}}^{-1}(V),V)$.
To show ’ii $\Longrightarrow$ i’, assume $(F_{Y}\circ F_{Y|\mathbb{X}}^{-1}(V),V)\geq_{c}(F_{Y}\circ F_{Y^{\prime}|\mathbb{X}^{\prime}}^{-1}(V),V)$. This implies by (B.10) and (B.12) the marginal constraint $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$, and by Proposition B.2 and Lemma B.3 that $P(F_{Y|\mathbb{X}}^{-1}(V)\leq q_{Y}(v),V\leq t)\geq P(F_{Y^{\prime}|\mathbb{X}^{\prime}}^{-1}(V)\leq q_{Y^{\prime}}(v),V\leq t)$ for all $t\in\mathbb{R}$ and $\lambda$-almost all $v\in(0,1)$. Consider the stop-loss function $x\mapsto(x-t)_{+}=-\min\{x,t\}+x$. Then, for $\lambda$-almost all $v\in(0,1)$, we obtain
$\mathbb{E}\left(P(Y\leq q_{Y}(v)\mid\mathbb{X})-t\right)_{+}$ $=-\mathbb{E}\min\{F_{Y|\mathbb{X}}(q_{Y}(v)),t\}+\mathbb{E}(P(Y\leq q_{Y}(v)\mid\mathbb{X}))$
$=-P(F_{Y|\mathbb{X}}^{-1}(V)\leq q_{Y}(v),V\leq t)+P(Y\leq q_{Y}(v))$
$\leq-P(F_{Y^{\prime}|\mathbb{X}^{\prime}}^{-1}(V)\leq q_{Y^{\prime}}(v),V\leq t)+P(Y^{\prime}\leq q_{Y^{\prime}}(v))$
$=\cdots=\mathbb{E}\left(P(Y^{\prime}\leq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})-t\right)_{+}$
for all $t\in\mathbb{R}$, where the second equality is due to Lemma B.4 and the inequality involves the marginal constraint. Using the marginal constraint, we obtain from the characterization of the convex order by the stop-loss order (see [45, Theorem 1.5.3 and Theorem 1.5.7]) that $P(Y\leq q_{Y}(v)\mid\mathbb{X})\leq_{cx}P(Y^{\prime}\leq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})$ for $\lambda$-almost all $v\in(0,1)$.
To show ’i $\Longrightarrow$ iii’, assume that $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\,.$ Then we obtain
@@
$\begin{split}P(q_{Y;\mathbb{X}}^{\uparrow U}(V)\leq q_{Y}(v),U\leq u)&=\int_{0}^{u}P\left(q^{\uparrow U}_{Y;\mathbb{X}}(V)\leq q_{Y}(v)\mid U=s\right)\mathsf{\,d}s\\ &=\int_{0}^{u}P\left(q^{\uparrow s}_{Y;\mathbb{X}}(V)\leq q_{Y}(v)\right)\mathsf{\,d}s\\ &=\int_{0}^{u}P\left(V\leq F_{s}(q_{Y}(v))\right)\mathsf{\,d}s=\int_{0}^{u}(\eta_{Y|\mathbb{X}}^{v})^{*}(s)\mathsf{\,d}s\\ &\leq\int_{0}^{u}(\eta_{Y^{\prime}|\mathbb{X}^{\prime}}^{v})^{*}(s)\mathsf{\,d}s\\ &=\cdots=P(q_{Y^{\prime};\mathbb{X}^{\prime}}^{\uparrow U}(V)\leq q_{Y^{\prime}}(v),U\leq u)\end{split}$ (B.15)
@@
for $\lambda$-almost all $v\in(0,1)$ and for all $u\in[0,1]$ with equality for $u=1,$ where we use for the inequality the characterization of $\preccurlyeq_{ccx}$ by the Schur order; see Theorem 2.2. The first equality is due to the disintegration theorem, the second equality holds by independence of $U$ and $V.$ For the third equality, we apply Proposition A.1i using that $F_{s}$ is a distribution function with associated quantile function $q_{Y;\mathbb{X}}^{\uparrow s}$; see (13). The fourth equality follows with (12) noting that $V$ is uniform on $(0,1)$.
@@
Using again $q_{Y;\mathbb{X}}^{\uparrow U}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y$ and $q_{Y^{\prime};\mathbb{X}^{\prime}}^{\uparrow U}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$ and using the marginal constraint, we apply Lemma B.3 to prove that (B.15) implies
$P\big(F_{Y}(q_{Y;\mathbb{X}}^{\uparrow U}(V))\geq v,U\geq u\big)\leq P\big(F_{Y^{\prime}}(q_{Y^{\prime};\mathbb{X}^{\prime}}^{\uparrow U}(V))\geq v,U\geq u\big)$ (B.16)
for all $(u,v)\in(0,1)^{2}$. Now, statement iii follows from (B.16) by Proposition B.2 where we use the identity in (B.11).
@@
To prove ’iii $\Longrightarrow$ i’, assume that $(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)\leq_{c}(F_{Y^{\prime}}\circ q_{Y^{\prime};\mathbb{X}^{\prime}}^{\uparrow U}(V),U)$. Applying Proposition B.2, we obtain (B.11) and the inequality (B.16) for all $(u,v)\in[0,1]^{2}$. Using that (B.11) is equivalent to (B.12) which implies the marginal constraint, we obtain from (B.16) by Lemma B.3 that
$P\big(q_{Y;\mathbb{X}}^{\uparrow U}(V)\leq q_{Y}(v),U\leq u\big)\leq P\big(q_{Y^{\prime};\mathbb{X}^{\prime}}^{\uparrow U}(V)\leq q_{Y^{\prime}}(v),U\leq u\big)$ (B.17)
@@
for $\lambda$-almost all $v\in(0,1)$ and for all $u\in(0,1)$, where $U$ and $V$ are independent and uniform on $(0,1)$. By a similar reasoning as in (B.15), it follows that
$\int_{0}^{u}(\eta_{Y|\mathbb{X}}^{v})^{*}(s)\mathsf{\,d}s$ $=P\big(q_{Y;\mathbb{X}}^{\uparrow U}(V)\leq q_{Y}(v),U\leq u\big)$
$\leq P\big(q_{Y^{\prime};\mathbb{X}^{\prime}}^{\uparrow U}(V)\leq q_{Y^{\prime}}(v),U\leq u\big)=\int_{0}^{u}(\eta_{Y^{\prime}|\mathbb{X}^{\prime}}^{v})^{*}(s)\mathsf{\,d}s$